OpenStudy (anonymous):

For a four-question multiple choice pop quiz, a student ti totally unprepared and randomly guesses the answer to each question. If each question has four options, then the probability of selecting the correct answer for any given question is 1/4 or .25. What is the probability that the student answers 3 correct. Can you walk me though this problem?

OpenStudy (anonymous):

1/4 = probability of one. 1/4 x 1/4 = probability of two 1/4 x 1/4 x 1/4 = probability of three So, the probability of the student getting three questions right by guessing is 1/64 or 0.015625